(2+2)/2-1+3*5

(2+2)/2-1+3*5

Grade:10

2 Answers

Ravi
9 years ago
What is the problem you are facing in solving this expression.
It’s simple maths. Apply the principle of BODMAS(1.Bracket open 2. Divide 3. Multiply 4. Addition 5. Subtract) in the sequence to get the answer.
.......................................................................................................................................
Harsh Patodia IIT Roorkee
9 years ago
Hi Student

We will use the BODMAS rule
(2+2)/2 -1 +3*5= 4/2 -1 +3*5= 2-1+3*5= 2-1+15 = 1+15=16

….….….….…...….….….….….….….….….….….….….…...